# Normality Calibration in Semi-supervised Graph Anomaly Detection

### v1

@@ Abstract
Graph anomaly detection (GAD) has attracted growing interest for its crucial ability to uncover irregular patterns in broad applications. Semi-supervised GAD, which assumes a subset of annotated normal nodes available during training, is among the most widely explored application settings. However, the normality learned by existing semi-supervised GAD methods is limited to the labeled normal nodes, often inclining to overfitting the given patterns. These can lead to high detection errors, such as high false positives. To overcome this limitation, we propose GraphNC, a graph normality calibration framework that leverages both labeled and unlabeled data to calibrate the normality from a teacher model (a pre-trained semi-supervised GAD model) jointly in anomaly score and node representation spaces. GraphNC includes two main components, anomaly score distribution alignment (ScoreDA) and perturbation-based normality regularization (NormReg). ScoreDA optimizes the anomaly scores of our model by aligning them with the score distribution yielded by the teacher model. Due to accurate scores in most of the normal nodes and part of the anomaly nodes in the teacher model, the score alignment effectively pulls the anomaly scores of the normal and abnormal classes toward the two ends, resulting in more separable anomaly scores. Nevertheless, there are inaccurate scores from the teacher model. To mitigate the misleading by these scores, NormReg is designed to regularize the graph normality in the representation space, making the representations of normal nodes more compact by minimizing a perturbation-guided consistency loss solely on the labeled nodes. Through comprehensive experiments on six benchmark datasets, we show that, by jointly optimizing these two components, GraphNC can 1) consistently and substantially enhance the GAD performance of teacher models from different types of GAD methods and 2) achieve new state-of-the-art semi-supervised GAD performance. Code will be made available at https://github.com/mala-lab/GraphNC.
@@ 1 Introduction
Graph anomaly detection (GAD), aiming to identify irregular nodes in a graph, has gained increasing attention due to its critical role in real applications such as detection of spams in social networks and frauds in financial networks (Pang et al., 2021a; Ma et al., 2021; Qiao et al., 2025). Semi-supervised GAD, which leverages a subset of labeled normal nodes to learn normal patterns, has gained significant attention in real-world applications (Qiao et al., 2024). This GAD setting is practical in that normal nodes often account for the majority in a graph, and thus, it is significantly less costly to obtain compared to abnormal samples that are scarce and their occurrence typically involves high financial/reputation loss. However, compared to unsupervised GAD that is more popular, less studies have been done on semi-supervised GAD. Furthermore, existing semi-supervised GAD methods are primarily built on the limited labeled normal nodes, preventing them from learning the exact normal pattern, thus inclining to overfit the annotated normality (Ding et al., 2019; Wang et al., 2021; Ding et al., 2021; Fan et al., 2020; Qiao et al., 2024). As a result, these methods suffer from high detection errors—e.g., normal nodes that are dissimilar to the labeled normal nodes are detected as anomalies (false positives) and vice versa (false negatives), especially the former one—due to the large overlapped anomaly scores for the normal and abnormal nodes. These issues can be observed in the results of a recent state-of-the-art (SOTA) method GGAD (Qiao et al., 2024) in Figs. 1 (a) and (b). The same phenomenon can also be observed in other semi-supervised methods like data reconstruction and one-class classification-based methods (see App. C).
To address these issues, we propose GraphNC, a novel graph normality calibration framework that exploits both labeled and unlabeled data to calibrate the normality learned from a teacher model (i.e., a pre-trained semi-supervised GAD model) jointly in anomaly score and representation spaces.
@@
  • GraphNC is a flexible framework where different teacher models can be plugged and played, achieving consistently enhanced GAD improvement across three types of teacher models, including data reconstruction-based, one-class-based, and anomaly-generation-based approaches. Further, GraphNC gets better performance when the teacher model is stronger. This is verfied by our comprehensive experiments on six benchmark datasets.
@@ 2 Related Work
Graph Anomaly Detection. Existing GAD methods can be broadly categorized into unsupervised, semi-supervised, and fully supervised approaches (Ma et al., 2021; Qiao et al., 2025). Unsupervised methods typically assume that no labeled nodes are known, and they learn normal patterns through some proxy task designs such as reconstruction, one-class classification, and adversarial learning (Ding et al., 2019; Wang et al., 2021; Ding et al., 2021). Although this setting is widely used in anomaly detection, they overlook the abundance of normal samples, where normal nodes are less costly to obtain than the abnormal nodes. The fully supervised methods, on the other hand, rely on labeled normal and abnormal nodes and formulate the task as an imbalanced classification problem (Liu et al., 2021a; Tang et al., 2022; Gao et al., 2023a; b; Chen et al., 2024; Liu et al., 2025). These supervised methods rely on large labeled data, limiting their applicability in real-world scenarios, especially those where anomalies are very infrequent and very costly to collect; they may overfit labeled anomalies, failing to generalize to unseen anomalies (Wang et al., 2025). Semi-supervised methods are generally more practical for real-world applications, as only a subset of labeled normal nodes is required (Qiao et al., 2024; 2025). Only limited work has been done in this line. GGAD (Qiao et al., 2024) is a recent work that utilizes the labeled normal nodes to generate synthetic outliers that mimic real anomalies, and then trains a binary classifier on this data. It also shows that popular unsupervised methods, such as reconstruction, one-class classification, or adversarial learning-based methods, can be adapted to the semi-supervised settings by minimizing their loss on the normal data only.
@@
Semi-supervised Learning. Semi-supervised learning aim to leverage a small amount of labeled data together with a large amount of unlabeled data to train models, aiming to achieve high performance while reducing the cost of data annotation (Yang et al., 2022; Chen et al., 2022b). Representative semi-supervised learning approaches include pseudo-labeling and consistency-based methods. The pseudo-labeling with confidence thresholding is highly successful and widely-adopted, which is typically built on the smoothness assumption that neighboring data points in the feature space tend to exhibit similar labels (Grandvalet & Bengio, 2004; Chen et al., 2023). However, such a mechanism requires a quantity–quality trade-off, which undermines the learning process, since incorporating more unlabeled data helps increase data coverage but often introduces noisy or unreliable pseudo-labels (Pham et al., 2021; Kage et al., 2024). The consistency-based model enforces prediction consistency under stochastic noise or data augmentation (Laine & Aila, 2016). Recent studies such as RankMatch (Mai et al., 2024), InterLUDE (Huang et al., 2024), and SCHOOL (Mo et al., 2024), have been proposed in different data modalities, including visual images, text, and graphs, aiming to encourage stable predictions across views, thereby enhancing generalization and reducing reliance on labeled data (Gui et al., 2024). However, they are mainly focused on classification tasks, different from our anomaly detection task that has only one-class labels and unbounded distribution in the anomaly class.
@@ 4.2 Anomaly Score Distribution Alignment (ScoreDA)
Formally, let $\mathcal{F}_{\mathcal{T}}$ be the pre-trained teacher model, $\mathcal{Y}^{\mathcal{T}}=\mathcal{F}_{\mathcal{T}}{(\mathbf{X},\mathcal{G};\Theta)}$ be the output scores for the entire node set, where $\Theta$ are the trainable parameters, and $\mathcal{Y}^{\mathcal{T}}=\{y_{1}^{\mathcal{T}},y_{2}^{\mathcal{T}},...,y_{N}^{\mathcal{T}}\}$ be the anomaly score distribution of samples, then we leverage a student model $\mathcal{F}_{S}$ with parameter $\Phi$ which is implemented as the combination of GNN and MLP layers (i.e., $\mathcal{F}_{GNN}$ with parametwers $\Omega$ and $\mathcal{F}_{MLP}$ with parameters $\phi$ respectively) to align the score set $\mathcal{Y}^{\mathcal{T}}$. For the student model, let $\mathcal{H}^{S}$ be the output of the GNN student model, for a node $v_{i}$, the representation is then denoted by ${\bf{h}}_{i}^{\mathcal{S}}=\mathcal{F}_{GNN}({\bf{x}}_{i},\mathcal{G};\Omega)$. On top of that, we apply a MLP layer to obtain the predicted score, $y^{\mathcal{S}}_{i}=\mathcal{F}_{MLP}({\bf{h}}_{i}^{\mathcal{S}};\phi)$. Finally, we employ the MSE loss to minimize the discrepancy between the anomaly score from the student model $\mathcal{Y}^{\mathcal{S}}$ and $\mathcal{Y}^{\mathcal{T}}$ yielded by the teacher model, which is formulated as follows:
@@ 4.3 Perturbation-Guided Normality Regularization (NormReg)
${{\cal L}_{{NormReg}}}=-\frac{1}{{\left|{{{\cal V}_{l}}}\right|}}\sum\limits_{{v_{i}}\in{\cal V}_{l}}{\left\|{{{\bf{h}}_{i}^{\mathcal{S}}}-{{\widetilde{\bf{h}}}_{i}^{\mathcal{S}}}}\right\|_{2}^{2}},$ (3)
@@
The proof can be found in App. D. According to Theorem 1, minimizing $\mathcal{L}_{NormReg}$ complements $\mathcal{L}_{ScoreDA}$ in reducing the intra-class score variance of normal classes in the student model, which can also be observed when comparing the scores of the normal class in Fig. 1(e) to that in Fig. 1(d), lowering both FPR and FNR, especially the FPR.
@@ 4.4 Training and Inference
5 Experiments
Datasets. We evaluate the effectiveness of the GraphNC across six real-world GAD datasets drawn from diverse domains, including social networks: Reddit (Kumar et al., 2019)), online shopping co-review networks: Amazon (Dou et al., 2020) and YelpChi (Dou et al., 2020), a co-purchase network: Photo (Shchur et al., 2018), a collaboration network: Tolokers (McAuley et al., 2015), and financial networks: T-Finance (Tang et al., 2022). Additional description and statistical results are provided in App. A. Note that some unsupervised methods, such as CoLA (Liu et al., 2021b), GRADATE (Duan et al., 2023), and HUGE (Pan et al., 2025), cannot be adapted to semi-supervised setting due to their design more relying on fully unlabeled data, which are excluded from our comparison.
Competing Methods. The competing methods include reconstruction-based methods: DOMINANT (Ding et al., 2019) and AnomalyDAE (Fan et al., 2020), one-class classification: OCGNN (Wang et al., 2021), adversarial learning based methods: AEGIS (Ding et al., 2021) and GAAN (Chen et al., 2020), affinity maximization-based methods: TAM (Qiao & Pang, 2023), and anomaly generation-based method: GGAD (Qiao et al., 2024). A more detailed description of competing methods can be found in the App. B. Except GGAD that is specifically designed for the semi-supervised setting, the competing methods are primarily unsupervised methods, which are adapted to our semi-supervised setting following (Qiao et al., 2024): reconstruction is performed using only the normal nodes, the one-class center calculation is based on only the normal nodes, the adversarial discriminator is trained with the normal nodes, and the affinity maximization is also applied on the normal nodes solely.
@@ 5 Experiments
5.1 Main Experiments

### v2

@@ Abstract
Semi-supervised graph anomaly detection (GAD), which assumes a subset of annotated normal nodes available during training, is among the most widely explored applications. However, the normality learned by existing semi-supervised GAD methods is limited to the labeled normal nodes, often inclining to overfitting the given patterns, thereby leading to high detection errors, such as high false positives. To overcome this limitation, we propose GraphNC, a graph normality calibration framework that leverages both labeled and unlabeled data to calibrate the normality from a teacher (a pre-trained semi-supervised GAD model) jointly in anomaly score and representation spaces. GraphNC includes two main components, anomaly score distribution alignment (ScoreDA) and perturbation-based normality regularization (NormReg). ScoreDA optimizes the anomaly scores of our model by aligning them with the score distribution yielded by the teacher. Due to accurate scores in most of the normal nodes and part of the anomaly nodes in the teacher, the alignment effectively pulls the anomaly scores of the two classes toward the two ends, resulting in more separable anomaly scores. To mitigate the misleading by inaccurate scores from the teacher, NormReg is designed to regularize the normality in representation space, making the representations of normal nodes more compact by minimizing a perturbation-guided consistency loss solely on the labeled nodes. Comprehensive experiments on six benchmarks demonstrate that GraphNC (1) consistently and substantially enhances the performance of teacher models from different GAD methods (2) achieves new state-of-the-art performance. Code will be made available at https://github.com/mala-lab/GraphNC.
Machine Learning, ICML
@@ 1 Introduction
Graph anomaly detection (GAD), aiming to identify irregular nodes in a graph, has gained increasing attention due to its critical role in real applications such as the detection of spams in social networks and frauds in financial networks (Pang et al., 2021a; Ma et al., 2021; Qiao et al., 2025). Semi-supervised GAD, which leverages a subset of labeled normal nodes to learn normal patterns, has gained significant attention in real-world applications (Qiao et al., 2024). This GAD setting is practical in that normal nodes often account for the majority in a graph, and thus, it is significantly less costly to obtain compared to abnormal samples that are scarce and their occurrence typically involves high financial/reputation loss. However, compared to unsupervised GAD that is more popular, less studies have been done on semi-supervised GAD. Furthermore, existing semi-supervised GAD methods are primarily built on the limited labeled normal nodes, preventing them from learning the exact normal pattern, thus inclining to overfit the annotated normality (Ding et al., 2019; Wang et al., 2021; Ding et al., 2021; Fan et al., 2020; Qiao et al., 2024). As a result, these methods suffer from high detection errors—e.g. , normal nodes that are dissimilar to the labeled normal nodes are detected as anomalies (false positives) and vice versa (false negatives), especially the former one—due to the large overlapped anomaly scores for the normal and abnormal nodes. These issues can be observed in the results of a recent state-of-the-art (SOTA) method GGAD (Qiao et al., 2024) in Figs. 1 (a) and (b). The same phenomenon can also be observed in other semi-supervised methods like data reconstruction and one-class classification-based methods (see App. C). To address these issues, we propose GraphNC, a novel graph normality calibration framework that exploits both labeled and unlabeled data to calibrate the normality learned from a teacher model (i.e. , a pre-trained semi-supervised GAD model) jointly in anomaly score and representation spaces.
@@
  • GraphNC is a flexible framework where different teacher models can be plugged and played, achieving consistently enhanced GAD improvement across three types of teacher models, including data reconstruction-based, one-class-based, and anomaly-generation-based approaches. Further, GraphNC gets better performance when the teacher model is stronger. This is verified by our comprehensive experiments on six benchmark datasets.
@@ 2 Related Work
Graph Anomaly Detection. Existing GAD methods can be broadly categorized into unsupervised, semi-supervised, and fully supervised approaches (Ma et al., 2021; Qiao et al., 2025). Unsupervised methods typically assume that no labeled nodes are known, and they learn normal patterns through some proxy task designs such as reconstruction, one-class classification, and adversarial learning (Ding et al., 2019; Wang et al., 2021; Ding et al., 2021). Although this setting is widely used in anomaly detection, they overlook the abundance of normal samples, where normal nodes are less costly to obtain than the abnormal nodes. The fully supervised methods, on the other hand, rely on labeled normal and abnormal nodes and formulate the task as an imbalanced classification problem (Liu et al., 2021a; Tang et al., 2022; Gao et al., 2023a, b; Chen et al., 2024; Liu et al., 2025). These supervised methods rely on large labeled data, limiting their applicability in real-world scenarios, especially those where anomalies are very infrequent and very costly to collect; they may overfit labeled anomalies, failing to generalize to unseen anomalies (Wang et al., 2025). Semi-supervised methods are generally more practical for real-world applications, as only a subset of labeled normal nodes is required (Zhou et al., 2022; Qiao et al., 2024, 2025). Only limited work has been done in this line. GGAD (Qiao et al., 2024) is a recent work that utilizes the labeled normal nodes to generate synthetic outliers that mimic real anomalies, and then trains a binary classifier on this data. It also shows that popular unsupervised methods, such as reconstruction, one-class classification, or adversarial learning-based methods, can be adapted to the semi-supervised settings by minimizing their loss on the normal data only.
@@
Semi-supervised Learning. Semi-supervised learning aim to leverage a small amount of labeled data together with a large amount of unlabeled data to train models, aiming to achieve high performance while reducing the cost of data annotation (Yang et al., 2022; Chen et al., 2022b). Representative semi-supervised learning approaches include pseudo-labeling and consistency-based methods. The pseudo-labeling with confidence thresholding is highly successful and widely-adopted, which is typically built on the smoothness assumption that neighboring data points in the feature space tend to exhibit similar labels (Grandvalet and Bengio, 2004; Chen et al., 2023). However, such a mechanism requires a quantity–quality trade-off, which undermines the learning process, since incorporating more unlabeled data helps increase data coverage but often introduces noisy or unreliable pseudo-labels (Pham et al., 2021; Kage et al., 2024). The consistency-based model enforces prediction consistency under stochastic noise or data augmentation (Laine and Aila, 2016). Recent studies such as RankMatch (Mai et al., 2024), InterLUDE (Huang et al., 2024), and SCHOOL (Mo et al., 2024), have been proposed in different data modalities, including visual images, text, and graphs, aiming to encourage stable predictions across views, thereby enhancing generalization and reducing reliance on labeled data (Gui et al., 2024). However, they are mainly focused on classification tasks, different from our anomaly detection task that has only one-class labels and unbounded distribution in the anomaly class.
@@ 4.2 Anomaly Score Distribution Alignment (ScoreDA)
Formally, let $\mathcal{F}_{\mathcal{T}}$ be the pre-trained teacher model, $\mathcal{Y}^{\mathcal{T}}=\mathcal{F}_{\mathcal{T}}{(\mathbf{X},\mathcal{G};\Theta)}$ be the output scores for the entire node set, where $\Theta$ are the trainable parameters, and $\mathcal{Y}^{\mathcal{T}}=\{y_{1}^{\mathcal{T}},y_{2}^{\mathcal{T}},...,y_{N}^{\mathcal{T}}\}$ be the anomaly score distribution of samples, then we leverage a student model $\mathcal{F}_{S}$ with parameter $\Phi$ which is implemented as the combination of GNN and MLP layers (i.e. , $\mathcal{F}_{GNN}$ with parameters $\Omega$ and $\mathcal{F}_{MLP}$ with parameters $\phi$ respectively) to align the score set $\mathcal{Y}^{\mathcal{T}}$. For the student model, let $\mathcal{H}^{S}$ be the output of the GNN student model, for a node $v_{i}$, the representation is then denoted by ${\bf{h}}_{i}^{\mathcal{S}}=\mathcal{F}_{GNN}({\bf{x}}_{i},\mathcal{G};\Omega)$. On top of that, we apply a MLP layer to obtain the predicted score, $y^{\mathcal{S}}_{i}=\mathcal{F}_{MLP}({\bf{h}}_{i}^{\mathcal{S}};\phi)$. Finally, we employ the MSE loss to minimize the discrepancy between the anomaly score from the student model $\mathcal{Y}^{\mathcal{S}}$ and $\mathcal{Y}^{\mathcal{T}}$ yielded by the teacher model, which is formulated as follows:
@@ 4.3 Perturbation-Guided Normality Regularization (NormReg)
${{\cal L}_{{NormReg}}}=\frac{1}{{\left|{{{\cal V}_{l}}}\right|}}\sum\limits_{{v_{i}}\in{\cal V}_{l}}{\left\|{{{\bf{h}}_{i}^{\mathcal{S}}}-{{\widetilde{\bf{h}}}_{i}^{\mathcal{S}}}}\right\|_{2}^{2}},$ (3)
@@
The proof can be found in App. D. According to Theorem 4.1, minimizing $\mathcal{L}_{NormReg}$ complements $\mathcal{L}_{ScoreDA}$ in reducing the intra-class score variance of normal classes in the student model, which can also be observed when comparing the scores of the normal class in Fig. 1(e) to that in Fig. 1(d), lowering both FPR and FNR, especially the FPR.
@@ 4.4 Training and Inference
5 Experiments
Datasets. We evaluate the effectiveness of the GraphNC across six real-world GAD datasets drawn from diverse domains, including social networks: Reddit (Kumar et al., 2019), online shopping co-review networks: Amazon (Dou et al., 2020) and YelpChi (Dou et al., 2020), a co-purchase network: Photo (Shchur et al., 2018), a collaboration network: Tolokers (McAuley et al., 2015), and financial networks: T-Finance (Tang et al., 2022). Additional description and statistical results are provided in App. A.
Competing Methods. The competing methods include reconstruction-based methods: DOMINANT (Ding et al., 2019), GADNR (Roy et al., 2024), ADA-GAD (He et al., 2024) and AnomalyDAE (Fan et al., 2020), one-class classification: OCGNN (Wang et al., 2021), adversarial learning based methods: AEGIS (Ding et al., 2021) and GAAN (Chen et al., 2020), affinity maximization-based methods: TAM (Qiao and Pang, 2023), and anomaly generation-based method: GGAD (Qiao et al., 2024). A more detailed description of competing methods can be found in the App. B. Except GGAD that is specifically designed for the semi-supervised setting, the competing methods are primarily unsupervised methods, which are adapted to our semi-supervised setting following (Qiao et al., 2024): reconstruction is performed using only the normal nodes, the one-class center calculation is based on only the normal nodes, the adversarial discriminator is trained with the normal nodes, and the affinity maximization is also applied on the normal nodes solely. Note that some unsupervised methods, such as CoLA (Liu et al., 2021b), GRADATE (Duan et al., 2023), and HUGE (Pan et al., 2025), cannot be adapted to semi-supervised setting due to their design more relying on fully unlabeled data, which are excluded from our comparison.
@@ 5 Experiments
Method Dataset
Amazon T-Finance Tolokers
AUROC
OT 0.9443 0.8228 0.5340
OT+NormReg 0.8956 0.8455 0.5843
OT+NormReg-Finetune 0.9060 0.8008 0.5396
OT+ScoreDA 0.9511 0.8300 0.5811
OT+ScoreDA+NormReg* 0.9452 0.8278 0.6492
OT+ScoreDA+NormReg 0.9613 0.8340 0.6505
AUPRC
OT 0.7922 0.1825 0.2502
OT+NormReg 0.6568 0.2413 0.2688
OT+NormReg-Finetune 0.7008 0.1406 0.2465
OT+ScoreDA 0.8189 0.2566 0.2567
OT+ScoreDA+NormReg* 0.8180 0.1971 0.3034
OT+ScoreDA+NormReg 0.8403 0.3667 0.3082
5.1 Main Experiments
@@ 5.2 GraphNC Enabled Three Teacher Models
5.3 Ablation Study
In this section, we perform an ablation study to evaluate the contribution of each component in GraphNC. To this end, several variants of GraphNC are introduced. (1) Only Teacher (OT) directly uses the pre-trained semi-supervised teacher model to derive the anomaly score. (2) OT+NormReg incorporates NormReg into the pre-training of the teacher model. (3) OT+NormReg+Finetune leverages NormReg to fine-tune the pre-trained teacher. (4) OT+ScoreDA includes the student model and adds ScoreDA on top of OT (i.e. , optimizing GraphNC using ScoreDA only). (5) OT+ScoreDA+NormReg* applies NormReg on top of OT+ScoreDA but it applies the NormReg loss on all nodes instead of the labeled normal nodes only. OT+ScoreDA+NormReg is the default full model using both ScoreDA and NormReg. The experimental results are shown in Table 3. The ablation study experiments on all datasets are provided in the App. C.2. We observe that OT+ScoreDA enhances the performance of OT, achieving the second-best results on most datasets in terms of AUPRC. This demonstrates that the ScoreDA component can effectively calibrate the normality learned from the teacher, ensuring better separation between normal and abnormal classes. However, OT+ScoreDA underperforms the full model (the last row) on all datasets, highlighting that the negative impact of inaccurate anomaly scores from the teacher model can be effectively mitigated by the NormReg component. On the other hand, applying NormReg directly to the teacher via either joint optimization (OT+NormReg) or finetuning (OT+NormReg+Finetune) can improve OT to some extent. This justifies the effectiveness of NormReg from a different perspective, but NormReg works best when working in the student model and combining with ScoreDA due to its joint effects in minimizing the detection errors discussed in our theoretical analysis. Besides, we also conduct the ablation study on other teacher models to further analyze the effectiveness of each module in GraphNC. Meanwhile, we evaluate the performance of GraphNC when the teacher performs poorly, the average performance improvement of GraphNC w.r.t. different accuracy levels, and the comparison of the number of true anomalous instances that are accurately detected for GGAD and GraphNC. All the results can be found in the App. C.
5.4 Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of GraphNC w.r.t. the loss moderator $\alpha$, mask ratio $\omega$, and training size $R$. A detailed analysis on time complexity can be found App. E.1.
Performance w.r.t. the training size $R$. To explore the impact of training size $R$, we compare GraphNC with four semi-supervised GAD methods using varying numbers of training normal nodes, as shown in Figs.4 (e-f). Similar with other competing methods, with the increase of training size $R$, our model GraphNC maintains the superiority over other competing methods, demonstrating the effectiveness of GraphNC using different scales of labeled training data.
Conclusion
In this paper, we propose GraphNC, a generic graph normality calibration framework designed to enhance existing semi-supervised GAD methods, achieving substantial and consistent performance improvements across different types of existing approaches. GraphNC is composed of two key components: ScoreDA and NormReg. The ScoreDA module aligns the anomaly scores with the anomaly score distribution produced by the teacher model, effectively pushing the anomaly scores for the normal and abnormal classes toward opposite ends. This alignment ensures a clearer separation between normal and abnormal nodes in the score space, thereby leading to a lower FPR and FNR. The NormReg module is introduced to reduce the impact of inaccurate scores in the anomaly score distribution served by the teacher, enabling GraphNC to learn more compact normal representations. GraphNC is comprehensively verified by empirical results on six GAD datasets, which show that it consistently enhances performance of three different types of teacher models, and it achieves the best performance.
Limitation and Future Work. Our empirical results suggest that GraphNC can still yield consistent improvements even when the teacher performs relatively poorly, because of the role of ScoreDA and NormReg. Nevertheless, the effectiveness of ScoreDA still depends on the teacher being able to offer at least coarse supervisory information. Future work will explore other supplementary supervision or expert knowledge to reduce its reliance on the teacher model.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Detailed Description of Datasets
  • YelpChi (Dou et al., 2020): Similar to Amazon-all, YelpChi-all incorporates three types of edges: R-U-R (reviews written by the same user), R-S-R (reviews of the same product with identical star ratings), and R-T-R (reviews of the same product posted within the same month). YelpChi-all is then constructed by merging these different relations into a single unified relation (Chen et al., 2022a; Qiao and Pang, 2023).
Appendix B Description of Baselines
  • GAAN (Chen et al., 2020): It leverages a generative adversarial network where fake graph nodes are generated. Then the covariance matrix for real nodes and fake nodes are computed to enhace the node. Finally, a discriminator is trained to recognize whether two connected nodes are from a real or fake node.
  • TAM (Qiao and Pang, 2023): It first reveals the ’one class homophily’ and introduces the new anomaly measure score ’local node affinity’. As the local node affinity can not be directly obtained based on the embedding, it is optimized on truncated graphs where non-homophily edges are removed iteratively. The local node affinity is calculated on the learned representations on the truncated graphs.
  • GADNR (Roy et al., 2024): It employs a GNN-based autoencoder to reconstruct each node’s attributes, degree, and Gaussian-approximated neighbor feature distribution, and uses the weighted sum of these reconstruction losses as the anomaly score.
  • ADA-GAD (He et al., 2024): It adopts a two-stage anomaly-denoised autoencoder framework that first pretrains on anomaly-reduced augmented graphs and then retrains decoders on the original graph, taking the final reconstruction error with a distribution regularization term as the anomaly score.
Appendix C Additional Experimental Results
Table 4: Key statistics of GAD datasets.
Datasets Amazon T-Finance Reddit YelpChi Tolokers Photo
#Nodes 11,944 39,357 10,984 45,941 11,758 7,484
#Edges 4,398,392 21,222,543 168,016 3,846,979 519,000 119,043
#Attributes 25 10 64 32 10 745
Anomaly 9.5% 4.6% 3.3% 14.52% 21.8% 4.9%
C.1 Score distribution visualization on other teacher model
In this section, we visualized the score distributions of the reconstruction-based method DOMINANT and the one-class classification method OCGNN, along with their corresponding GraphNC-enabled models on the Amazon. As shown in the Fig. 5 & 6, we observe that the GraphNC-enabled model achieves better separation between normal and abnormal classes, further demonstrating that GraphNC can substantially and consistently enhance performance.
(a) DOMINANT
(b) GraphNC(DOMINANT)
Figure 5: The score distribution of DOMINANT along the corresponding GraphNC enabled DOMINANT on Amazon (Dou et al., 2020).
(a) OCGNN
(b) GraphNC(OCGNN)
Figure 6: The score distribution of OCGNN along the corresponding GraphNC enabled OCGNN on Amazon (Dou et al., 2020).